OpenStudy (anonymous):

|y-8|+7=3 Step by step please.

OpenStudy (anonymous):

y+8+7=3 ( absolute value means no negatives) y+15=3 subtract 15 on both sides y=-12

OpenStudy (anonymous):

You can add them even though they have the | | ?

OpenStudy (blacksteel):

amk's answer is wrong. The absolute value sign means that the term inside the sign is made positive. So: |y-8| + 7 = 3. |y-8| = -4 Since |y-8| must be at least 0, this equation has no solution.

OpenStudy (anonymous):

Thanks. :)

OpenStudy (anonymous):

wait i made a mistake my bad i dont know what i was thinking

OpenStudy (anonymous):

It's fine.